# Bloom: Designing for LLM-Augmented Behavior Change Interactions

### v1

@@ 6.2.4. User Experience
6.3. Wearable Data
We now turn to objective measures of PA outcomes, complementing the self-reported survey outcomes. Table 7 presents coefficient means, standard errors, and significance level for each hypothesis from our mixed‑effects model. Across all four outcomes—step count, active energy burned (kcal), exercise time (min), and distance walking/running (km)—mean PA during the 4‑week study period was significantly higher than mean baseline PA (H1), corresponding to moderate-to-large standardized effects (Cohen’s d=0.59d=0.59-$0.75$). On average, participants walked 1,680 more steps/day ($p<0.001$, $d=0.75$), burned 87.1 more kcal/day ($p<0.01$, $d=0.59$), spent 13.2 more minutes/day exercising ($p<0.01$, $d=0.60$), and walked 0.76 km/day more ($p<0.001$, $d=0.74$) during the study period than at baseline. These effects translate to substantial increases in activity, doubling the proportion of participants meeting the recommended 150 min/week of exercise (for Disease Control and Prevention, 2022), increasing from 36% at baseline to 72% during the study (41% $\rightarrow$ 71% in control; 31% $\rightarrow$ 74% in treatment).
We did not find significant treatment–control differences in either the overall daily activity change from baseline to the four-week study period ($p=0.78$ to $1.0$) or the weekly rate of change ($p=0.77$ to $1.0$). However, the results showed consistent directional patterns. For the overall change from baseline to the study period (H2), every outcome had a negative coefficient, indicating that from baseline to the study period, the treatment group experienced somewhat smaller, though not significantly different, average gains than the control group. Regarding weekly persistence (H3), the treatment coefficients are uniformly positive, indicating activity in the treatment condition declined less (or even increased) per study week relative to control.
In summary, H1 was confirmed: both groups showed large and significant increases in PA over the baseline. H2 and H3 were not statistically supported, though consistent directional patterns suggest that the control group achieved slightly larger initial increases, while the treatment group maintained activity levels more consistently across weeks. However, we caution that the present sample was underpowered to detect small differences using statistical testing, and interpretations should be viewed as hypothesis generating rather than confirmatory.
@@ 6.4. Plan Data
Next, we examine the quality and personalization of participants’ weekly PA plans (Figure 8). First, we assess plan balance by computing a composite score measuring the number of different exercise categories included in each user’s weekly plan (+1 for each of aerobic, strength, flexibility; total score out of three), following the American College of Sports Medicine guidelines (of Sports Medicine, 2025). Treatment participants had more balanced plans, with a mean (SD) of 1.90 (0.54) categories per weekly plan, compared to 1.43 (0.55) in the control group.
Second, we measure the number of unique activity types per plan. The treatment group exhibited greater variety, with a mean (SD) of 3.77 (1.40) different activity types compared to 2.51 (1.20) in control. Next, we evaluated average plan completion rates (duration of completed activities/duration of all planned activities) and observed a slightly higher completion rate of 87% (14) in the treatment condition compared to control with 83% (21). Finally, we analyze the number of plan edits made by each participant per week. Treatment participants made 6.88 (7.27) plan edits on average, while control made 5.57 (3.97) edits. Together, these results suggest higher plan quality and a higher degree of personalization and engagement in the treatment condition.
@@ 7. Discussion
Our study provides rich insights into how integrating an LLM coaching chatbot with established behavior change interactions can impact PA behaviors and associated psychological outcomes. Bloom’s LLM-augmented interventions led to positive changes in participants’ mindsets, supported more flexible and personalized planning, and substantially increased app usage duration. In this section, we discuss implications for designing LLM-augmented behavior change interventions, the roles of LLMs in behavioral health systems, and limitations to address in future research.
@@ 7.1. The Role of LLMs in Behavioral Health Support: Coaching and Personalization
In Bloom, the LLM served two complementary roles: (1) as a conversational coaching agent that engaged in supportive conversations, and (2) as a personalization mechanism for tailoring interventions based on qualitative context. Whereas most prior work on LLMs in behavioral health has conceptualized their potential primarily in terms of personalization (e.g., customizing plans, optimizing just-in-time nudges, or extracting insights from data), our findings highlight that LLMs’ role as a relational agent providing motivational and non-judgmental support was more impactful for participants in early stages of behavior change and low baseline levels of activity.
@@
In our interviews, participants more often attributed their improved mindsets toward PA—increased enjoyment, greater identification with being an active person, greater self-confidence and self-compassion, and nuanced perspectives on what counts as exercise—to Beebo’s collaborative and supportive tone. These qualitative insights are consistent with our survey findings, in which treatment participants experienced larger pre-post gains in PA adequacy mindset and satisfaction with PA. Although participants appreciated that the LLM’s support was personalized to their unique circumstances, they more often attributed their ability to stick to their weekly plans to the encouragement and flexibility provided by the LLM along the way (e.g., rescheduling missed activities or suggesting alternatives), rather than to the personalized plans themselves. This is mirrored in our quantitative findings, where treatment participants demonstrated greater plan variety and adherence without a corresponding increase in objective PA outcomes.
@@
These findings align with existing behavior change theory, including principles of MI (Miller and Rollnick, 2023) and the TTM (Prochaska and Velicer, 1997), both of which emphasize that changing thoughts, feelings, and attitudes toward a behavior is crucial for individuals who are just starting out. Shifts in mindset (i.e., positive beliefs about the benefits of exercise) carry benefits to both psychological and physiological health (Crum and Langer, 2007) independent of immediate behavior change. Moreover, these changes in mindset may also represent shifts in decisional balance (i.e., stronger perceived pros and fewer perceived cons of PA), which is predictive of sustainable, longer-term behavior change in the TTM (Prochaska and Velicer, 1997). Thus, while immediate quantitative outcomes were not significantly different, the impact of LLM coaching on mindsets could potentially yield measurable behavioral changes over an extended timeframe, warranting longer-term studies.
@@
However, it remains an open question whether more granular LLM-driven personalization or more prescriptive coaching agents might better serve individuals who are already highly active or in later stages of change. These individuals may require more specific, performance-oriented support. Nonetheless, populations in early stages of change and/or with low baseline activity levels are particularly important to support as they stand to benefit most from behavioral health interventions.
7.2. Leveraging Qualitative Context to Foster Agency
While participants rarely identified Beebo as the single most useful feature, they described its role in supporting and reinforcing accountability provided by other app features, such as plans and notifications. Across our interviews, participants consistently valued the way Beebo’s interactions helped them feel a stronger sense of agency over their PA: agency in setting their own goals, adjusting plans as barriers arose, and reframing everyday activities as legitimate forms of exercise.
Participants highlighted Beebo’s role in collaborative goal-setting and appreciated how Beebo asked to reschedule missed activities rather than abandon them, or offered to substitute alternatives that better matched their schedules or abilities. Participants enjoyed notifications from Beebo that encouraged reflection on progress and motivation, which were perceived not as tasks to be completed but reminders of their own goals and motivations. Survey results bolster these findings, as treatment participants reported greater shared decision-making around goal-setting and perceived greater insight into their personal data.
@@ 7.2. Leveraging Qualitative Context to Foster Agency
Many of the interactions participants valued most were only possible because Beebo could leverage qualitative context expressed in natural language. For example, collaborative goal-setting requires that the agent asks about and adapts to personal constraints (e.g., childcare or social commitments) or physical abilities. Notifications that remind participants of their long-term goals and motivations necessitate knowledge of these goals and motivations. Others described how Beebo helped reframe activities they already enjoyed (e.g., gardening or walking at the mall) as legitimate forms of exercise, which reduced the perceived burden of “exercising” and made PA feel more achievable.
These findings highlight a design opportunity for LLM-augmented interactions to enhance agency not simply by personalizing content, but by contextualizing support in ways that help participants recognize and build upon their own capacities. Promoting agency is also strongly aligned with principles of health coaching and MI. For future systems, this could include integrating additional context sources (e.g., calendar or weather data) to proactively anticipate barriers, or reframing data visualizations around participants’ own goals to celebrate progress on their own terms rather than merely reporting quantitative metrics. Our results suggest that LLM augmentation is particularly useful for agency-enhancing interventions because qualitative context allows interventions to feel personally relevant and aligned with participants’ own motivations.
@@ 7.3. Social and Relational Cues: Opportunities and Risks for Promoting Engagement
Participants in the treatment condition exhibited increased engagement across all screens of the Bloom app. While multiple factors may have contributed to this, interview feedback suggests that the perceived social and relational nature of interactions with Beebo played an important role. Many participants explicitly said that chatting with Beebo felt like talking to a person, even while recognizing it was an AI. This mirrors findings from pre‑LLM health coaching systems (Bickmore and Picard, 2005; Mitchell et al., 2021), as well as recent LLM‑based coaching (Jörke et al., 2025), and aligns with long‑standing evidence that people respond socially to computers and media (Reeves and Nass, 1996). Participants identified several qualities that fostered this sense of presence, mentioning the bee avatar, Beebo’s empathetic and non-judgmental tone, proactive but optional messaging, and references back to prior conversations. Unlike the control condition, LLM-generated notifications and chatbot interactions were frequently described as interactions with an accountability partner rather than simple reminders.
@@
These social and relational cues carry both benefits and risks. On the one hand, they helped participants maintain accountability and fostered positive mindsets around exercise. One implication of this is that increasing anthropomorphic qualities by using more personalized and empathetic messaging, human-like avatars, or expanding conversational topics beyond strictly health-related content could strengthen engagement and accountability. On the other hand, such designs may foster emotional attachment or over‑reliance on the chatbot, raising concerns about self‑efficacy and intrinsic motivation. Prior work cautions that support from chatbots can differ in important and controversial ways from support provided by humans (Pan and De Graaf, 2025) and recent studies point to the mental health risks associated with affective use of LLM chatbots (McCain et al., 2025; Phang et al., 2025). Conversely, interfaces that minimize anthropomorphism or rely solely on UI-based, non-conversational interactions (e.g., simple free-response input) may mitigate these risks but may lose out on the motivational and accountability benefits we observed.
@@
Importantly, our findings show that social and relational cues should not be seen as the only means to foster engagement. For instance, Bloom’s ambient display was perceived as playful and appealing, was highly popular, and contributed to engagement and motivation. Prior work on narrative for behavior change (Murnane et al., 2020) suggests that incorporating narrative structures, where an LLM avatar could serve as a character in a story or interactive guide, could further enhance engagement and motivation without relying on increased anthropomorphism. Additionally, integrating LLM-augmented interactions within systems that facilitate human social support, where LLMs might mediate or support peer-to-peer interactions, offers another promising approach to foster accountability and engagement (Mitchell et al., 2021).
@@ 7.4. Design Challenges and Future Directions
While Bloom received largely positive feedback, participants identified several areas for improvement. For example, Beebo’s verbosity and repetitiveness were common complaints. In preliminary experiments, shortening responses via prompting resulted in reduced perceived empathy and support. We decided to allow our agent to produce longer responses based on this pilot feedback, even though verbosity is associated with lower-quality counseling (Pérez-Rosas et al., 2019). Addressing verbosity concerns will likely require larger scale data collection and fine-tuning beyond the prompting approaches used in this study. Meanwhile, our notification generation prompt included the last ten notifications along with a diversity prompt, yet they were still perceived as repetitive by some participants. Future iterations could incorporate more sophisticated memory systems to minimize repetitive or overly generic responses.
@@
Similarly, interactions sometimes felt rigid or ”bot-like,” potentially stemming from our dialogue state chain enforcing hard bounds on the conversation topic. While removing the dialogue state chain allowed the agent to more naturally adapt to a conversation, it also allowed it to veer off course and lose track of the conversation’s goals, mirroring prior findings in multi-turn settings (Laban et al., 2025). Future research on goal-directed, multi-turn conversations could improve LLMs’ conversational flexibility while ensuring they achieve conversational goals and stay within fixed conversational bounds.
@@
In addition, we encountered several issues attributable to inconsistent tool use. Participants occasionally experienced frustrating interactions due to hallucinated or inaccurate tool calls, particularly around scheduling and activity tracking. For example, hallucinated tool calls (e.g., extraneous or incorrect arguments) could lead to errors in plan edits, upon which the agent would tell the user they would “fix the issue later,” even though it did not have this capability. Moreover, health data fetching was rarely proactively initiated by the model, mirroring findings from GPTCoach (Jörke et al., 2025). Improving the robustness of tool use will likely also necessitate fine-tuning.
@@
Finally, several usability challenges emerged from inherent challenges in integrating conversational agents with complex user interfaces. The open-ended chat interface and Beebo’s conversational flexibility surfaced several false affordances: participants frequently posed questions about app features or bugs directly to Beebo (e.g., why workouts were not syncing, summaries not loading, and check-in not starting), which Beebo frequently did not have the knowledge or capabilities to address. Adding to their frustration, the agent would sometimes hallucinate an incorrect answer instead of stating that it did not know. These results might also explain the larger variance in post-study usability scores in the treatment condition. Addressing these concerns requires both clear communication of agent capabilities, either in the UI or chat, to calibrate participants’ mental models as well as preventing problematic model hallucinations.

### v2

@@ 6.3. Wearable Data
Plan quality and personalization by condition. Four scatter plots compare treatment (green) and control (gray) groups. Plot A shows plan balance, with treatment participants including more activity categories per plan. Plot B shows unique activity types, where treatment participants created plans with more unique types. Plot C shows plan completion rate, which was slightly higher in the treatment group. Plot D shows plan edits, with treatment participants making more weekly edits than controls.
@@
Average daily application usage. Two scatter plots compare treatment (green) and control (gray) groups. Plot A shows total daily app usage, with treatment participants spending over five times more time in the app. Plot B shows app usage by screen, where treatment participants used every screen (Today, Insights, Plan, and Chat) more than control.
@@ 6.4. Plan Data
Next, we examine the quality and personalization of participants’ weekly PA plans (Figure 8). First, we assess plan balance by computing a composite score measuring the number of different exercise categories included in each user’s weekly plan (+1 for each of aerobic, strength, flexibility; total score out of three), following the ACSM guidelines (of Sports Medicine, 2025). Treatment participants had more balanced plans, with a mean (SD) of 1.9 (0.5) categories per weekly plan, compared to 1.4 (0.6) in the control group.
Second, we measure the number of unique activity types per plan. The treatment group exhibited greater variety, with a mean (SD) of 3.8 (1.4) different activity types compared to 2.5 (1.2) in control. Next, we evaluated average plan completion rates (duration of completed activities/duration of all planned activities) and observed a slightly higher completion rate of 87% (14) in the treatment condition compared to control with 83% (21). Finally, we analyze the number of plan edits made by each participant per week. Treatment participants made 6.9 (7.3) plan edits on average, while control made 5.6 (4.0) edits. Together, these results suggest higher plan quality and higher degree of personalization and engagement in the treatment condition.
@@ 7. Discussion
Code Description
Qualitative Coding
Q1 Participants in both conditions most frequently described plans, ambient displays, and notifications—not chat—as the main sources of accountability and drivers of PA.
Q2 Beebo reinforced the accountability provided by other features (e.g., plans) by supporting flexible adjustments and giving personalized advice to overcome barriers.
Q3 Treatment participants described changes in their beliefs about PA and their own capabilities with greater specificity and nuance, frequently crediting Beebo’s supportive, empathetic, and non-judgmental tone.
Q4 Treatment participants described Beebo’s support in relational terms, referring to notifications as “coming from” Beebo, characterizing Beebo as an accountability partner, or comparing its support to that of a human.
Q5 Participants suggested that Beebo could “push” them more, provide more specific advice, and reduce repetitive messaging.
Daily, Weekly, and Pre-Post Surveys
S1 Self-reported PA levels increased substantially in both groups with similar magnitude. Exercise stage of change advanced slightly more among treatment participants.
S2 Treatment participants showed larger gains in PA adequacy mindset, satisfaction with PA and physical health, identifying as an active person, hopefulness about their health, and perceptions that exercising is easy.
S3 Usability scores declined pre- to post-study in both groups, with larger and more variable declines in the treatment condition.
S4 Treatment participants rated user experience, advice quality, shared decision-making, insight into data, and human-like support more highly overall, while user experience and advice quality declined pre- to post-study in both groups.
Wearable Data
W1 Mean PA levels during the study period were significantly higher than pre-study baseline levels (H1).
W2 Treatment-control differences during the study period (H2/H3) were statistically inconclusive but consistently showed the same directional patterns: the control group showed larger initial increases but declined more over time, while the treatment group showed smaller initial gains but maintained their PA levels more consistently.
Plan Data
P1 Treatment participants created more balanced and varied weekly plans with greater coverage of activity categories and more unique activity types.
P2 Plan completion rates were slightly higher in treatment (87%) than control (83%).
P3 Treatment participants made more weekly plan edits than control participants.
App Usage Data
A1 Treatment participants spent 5.6x as much daily time in the app on average and had higher usage time across all screens.
A2 App usage declined over time in both conditions but remained substantially higher in treatment across the study period.
Table 8. Summary of Findings. We report a summary of qualitative, survey, wearable, plan, and app usage findings, organized by codes that are referenced in Section 7.
In this section, we synthesize our findings into three themes, discussing the role of LLMs as a conversational coach versus a personalization mechanism, the use of qualitative context to promote agency, and the design trade-offs introduced by social and relational cues. Throughout, we reference findings using the codes summarized in Table 8. We conclude with key design challenges and opportunities for future improvement.
@@ 7.1. The Role of LLMs in Behavioral Health Support: Coaching and Personalization
In Bloom, the LLM served two complementary roles: (1) as a conversational coaching agent that engaged in supportive conversations, and (2) as a personalization mechanism for tailoring interventions based on qualitative context. Our findings suggest that these two roles had different impacts on participant experiences and behaviors. Plan data indicate that LLM-driven personalization produced higher quality and more personalized plans in the treatment group (P1-2), while user experience surveys show that treatment participants rated their support as more personalized and actionable (S4). However, these improvements did not translate into large differences in short-term PA compared to the no-LLM control (S1, W2). Instead, qualitative interviews indicate that the LLM’s role as a conversational coach had a greater influence on psychological and motivational outcomes associated with longer-term change (Q3).
@@
In our interviews, participants frequently attributed their improved mindsets toward PA—increased enjoyment, greater identification with being an active person, more self-confidence and self-compassion, and nuanced perspectives on what counts as exercise—to Beebo’s collaborative and supportive tone (Q3). Although they appreciated how Beebo personalized their plans, they more often attributed their ability to stick to their plans to Beebo’s encouragement and flexible support along the way, such as rescheduling missed activities or suggesting alternatives (Q2). Given that both conditions most commonly attributed their improved accountability and PA to plans, notifications, and the ambient display, not Beebo itself (Q1), Beebo’s effect in the treatment condition appeared to manifest primarily through improved mindsets. These qualitative insights align with our survey findings showing larger gains in mindset-related outcomes in the treatment condition (S2).
@@
These findings are consistent with behavior change theory, including principles of motivational interviewing (Miller and Rollnick, 2023) and the TTM (Prochaska and Velicer, 1997), which emphasize that changing thoughts, feelings, and attitudes toward a behavior are especially important for individuals in early stages of change. Shifts in mindset (i.e., positive beliefs about the benefits of exercise) carry benefits to both psychological and physiological health (Crum and Langer, 2007) independent of immediate behavior change. Moreover, changes in mindset may also represent shifts in decisional balance (i.e., stronger perceived pros and fewer perceived cons of PA), which is predictive of sustainable, longer-term change in the TTM (Prochaska and Velicer, 1997). Thus, even in the absence of immediate behavioral differences, the impact of LLM coaching on mindsets could potentially yield measurable behavioral changes over an extended timeframe, warranting longer-term studies.
@@
Whereas most prior work on LLMs in behavioral health has conceptualized their potential primarily in terms of personalization (e.g., customizing plans, optimizing just-in-time nudges, or extracting insights from data), our findings highlight that LLMs’ role as a motivational and non-judgmental coach was more impactful for participants in early stages of behavior change and low baseline levels of activity. It remains an open question whether more granular LLM-driven personalization or more prescriptive coaching agents might better serve individuals who are already highly active or in later stages of change, who may require more specific, performance-oriented support. Nonetheless, populations in early stages of change and/or with low activity levels are particularly important to support as they stand to benefit most from behavioral health interventions.
[…]
Implications for Design: For people in early stages of behavior change or with low PA levels, LLM-driven personalization alone may not meaningfully increase short-term behavior change beyond traditional interventions. Instead, consider pairing personalization with coaching to promote positive mindsets and foster engagement.
7.2. Leveraging Qualitative Context to Promote Agency
Participants consistently described Beebo as most helpful when it supported them in sticking to the plans they had set, as opposed to prescribing what they ought to do (Q2). While Beebo was rarely cited as the single most useful feature, a common theme across participants’ accounts was how Beebo’s flexible, non-prescriptive, and collaborative support strengthened their sense of agency over their PA. Many appreciated that Beebo asked for their opinion before offering advice, fostering a sense of control (Q3). They valued how Beebo asked to reschedule missed activities rather than abandon them, or offered to substitute alternatives that better matched their schedules or abilities (Q2). Notifications from Beebo were often described as reminders of their own intentions rather than obligations (Q4). Participants also reported how Beebo helped broaden their definition of activities that “count” as PA (Q3).
These findings align with prior work in HCI on goal setting and personalization. For instance, Lee et al. (2015) propose a reflective strategy in which participants answer reflective questions before setting a weekly exercise plan, finding that their Fitbit Plan system led to higher motivation and step count than automatically personalized plans. Niess & Woźniak’s (Niess and Woźniak, 2018) Tracker Goal Evolution Model describe how qualitative goals emerge from internal hedonic and eudaimonic needs, arguing that trackers should support users in translating abstract qualitative goals into concrete quantitative goals. Ekhtiar et al.’s (Ekhtiar et al., 2023) review of goal setting in personal informatics finds that systems should better support self-efficacy and reduce burden during goal setting. Saksono & Parker’s (Saksono and Parker, 2024) Socio-Cognitive Framework for HCI demonstrates that engaging with people’s aspirations increases positive outcome expectations, which in turn promote health behaviors. Promoting agency is also strongly aligned with principles of health coaching and motivational interviewing.
@@ 7.2. Leveraging Qualitative Context to Promote Agency
The agency-promoting interactions participants valued most fundamentally depended on the LLM’s ability to interpret and leverage qualitative context expressed in natural language. For example, effective collaborative goal setting requires an understanding of personal constraints such as childcare responsibilities, variable work schedules, or temporary injuries. Notifications that remind participants of their long-term goals and motivations necessitate knowledge of those motivations. Participants described how Beebo helped reframe activities they already enjoyed (e.g., gardening or walking at the mall) as legitimate forms of exercise, which depends on knowledge of their prior beliefs.
Our findings highlight a unique design opportunity for LLM-augmented interactions to leverage qualitative context to better promote users’ agency. Although pre-LLM systems could collect qualitative goals and experiences through free text input, their ability to tailor support was typically limited to fixed, rule-based adaptations, or depended on heavy user effort or human intermediaries. In contrast, LLMs’ ability to flexibly interpret qualitative context expressed in natural language enables non-prescriptive, non-judgmental interactions that center users’ own goals and aspirations. For example, future systems could support collaborative goal setting that engages more deeply with participants’ values and hold space for multiple complex, evolving, and potentially conflicting goals (Agapie et al., 2022). They could also aid in data collection and interpretation centered around participants’ own goals to support progress on their own terms, better support reflection that translates qualitative to quantitative goals (Niess and Woźniak, 2018), or integrate additional context sources (e.g., calendar or weather data) to proactively anticipate barriers.
[…]
Implications for Design: LLMs’ access to qualitative context affords support that more deeply engages with participants’ goals, barriers, and lived experiences. Consider applying this capability to design interactions that promote agency.
@@ 7.3. Social and Relational Cues: Opportunities and Risks for Promoting Engagement
Participants in the treatment condition exhibited increased engagement across all screens of the Bloom app (A1-2). While multiple factors may have contributed to this, interview feedback suggests that the perceived social and relational nature of interactions with Beebo played an important role (Q3-4). Many participants explicitly said that chatting with Beebo felt like talking to a person (Q4), even while recognizing it was an AI. This mirrors findings from pre-LLM health coaching systems (Bickmore and Picard, 2005; Mitchell et al., 2021), as well as recent LLM-based coaching (Jörke et al., 2025), and aligns with long-standing evidence that people respond socially to computers and media (Reeves and Nass, 1996). Participants identified several qualities that fostered this sense of presence, mentioning the bee avatar, Beebo’s empathetic and non-judgmental tone, proactive but optional messaging, and references back to prior conversations (Q3-4). Unlike the control condition, LLM-generated notifications and chatbot interactions were frequently described as interactions with an accountability partner rather than simple reminders (Q4). Interestingly, Bloom was able to engage with qualities of health informatics tools that Saksono and Parker (2024) describe as typically enabled by human social support, such as people’s long-term aspirations, sense of belongingness sustained through frequent meaningful interactions, and impediments to PA. This led to many of the same benefits as human social support, such as an increase in the perceived benefits of PA (Q3, S2).
@@
These social and relational cues carry both benefits and risks. On the one hand, they helped participants maintain accountability and fostered positive mindsets around exercise. One implication of this is that increasing anthropomorphic qualities by using more personalized and empathetic messaging, human-like avatars, or expanding conversational topics beyond strictly health-related content could strengthen engagement and accountability. On the other hand, such designs may foster emotional attachment or over-reliance on the chatbot, raising concerns about self-efficacy and intrinsic motivation. Prior work cautions that support from chatbots can differ in important and controversial ways from support provided by humans (Pan and De Graaf, 2025) and recent studies point to the mental health risks associated with affective use of LLM chatbots (McCain et al., 2025; Phang et al., 2025). Conversely, systems that minimize anthropomorphism or rely solely on UI-based, non-conversational interactions (e.g., free-response input) may mitigate these risks but may lose out on the motivational and accountability benefits we observed.
@@
Importantly, our findings show that social and relational cues should not be seen as the only means to foster engagement. For instance, Bloom’s ambient display was perceived as playful and appealing, was highly popular, and contributed to engagement and motivation (Q1). Prior work on narrative for behavior change (Murnane et al., 2020; Saksono et al., 2020) suggests that incorporating narrative structures, where an LLM avatar could serve as a character in a story or interactive guide, could further enhance engagement and motivation without relying on increased anthropomorphism. Additionally, integrating LLM-augmented interactions within systems that facilitate human social support (Saksono and Parker, 2024), where LLMs might mediate or support peer-to-peer interactions, offers another promising approach to foster accountability and engagement.
@@
[…]
Implications for Design: Relational cues are a powerful tool for promoting engagement and positive mindsets, but they should be applied with care. Consider pairing them with alternative engagement strategies, such as ambient displays, narrative, or human social support.
@@ 7.4. Design Challenges and Future Directions
While Bloom received largely positive feedback, participants identified several areas for improvement. For example, Beebo’s verbosity and repetitiveness were common complaints. In preliminary experiments, shortening responses via prompting resulted in reduced perceived empathy and support. We decided to allow our agent to produce longer responses based on this pilot feedback, even though verbosity is associated with lower-quality counseling (Pérez-Rosas et al., 2019). Addressing verbosity concerns will likely require larger scale data collection and training efforts beyond the prompting approaches used in this study. Meanwhile, our notification generation prompt included the last ten notifications along with a diversity prompt, yet they were still perceived as repetitive by some participants. Future iterations could incorporate more sophisticated memory systems or diversity sampling to minimize repetitive or generic responses.
@@
Similarly, interactions sometimes felt rigid or “bot-like,” potentially stemming from our dialogue state chain enforcing hard bounds on the conversation topic. More concerningly, the dialogue state chain placed Beebo in the role of steering conversations, potentially limiting participants’ agency. Many participants gradually shifted to short, one-word replies over time as questions became repetitive, limiting the context available to the agent. While removing the dialogue state chain allowed the agent to more naturally adapt to a conversation, it also allowed it to veer off course and lose track of the conversation’s goals, mirroring prior findings in multi-turn settings (Laban et al., 2025). Future research on goal-directed, multi-turn conversations could improve LLMs’ conversational flexibility to leave participants more space to steer the conversation while ensuring they achieve conversational goals and stay within fixed conversational bounds.
@@
In addition, we encountered several issues attributable to inconsistent tool use. Participants occasionally experienced frustrating interactions due to hallucinated or inaccurate tool calls, particularly around scheduling and activity tracking. For example, hallucinated tool calls (e.g., extraneous or incorrect arguments) could lead to errors in plan edits, upon which the agent would tell the user they would “fix the issue later,” even though it did not have this capability. Moreover, health data fetching was rarely proactively initiated by the model, mirroring findings from GPTCoach (Jörke et al., 2025). Improving the robustness of tool use may also necessitate multi-agent systems and/or finetuning.
@@
Finally, several usability challenges emerged from inherent challenges in integrating conversational agents with complex user interfaces. The open-ended chat interface and Beebo’s conversational flexibility surfaced false expectations about the system’s capabilities: participants frequently posed questions about app features or bugs directly to Beebo (e.g., why workouts were not syncing, summaries not loading, and check-in not starting), which Beebo frequently did not have the knowledge or capabilities to address. Adding to their frustration, the agent would sometimes hallucinate an incorrect answer instead of stating that it did not know. These results may explain the larger variance in post-study usability scores in the treatment condition. Addressing these concerns requires both clear communication of agent capabilities, either in the UI or chat, to calibrate participants’ mental models as well as preventing problematic model hallucinations.